Attendees: Bruce Glagola, George Srajer, Bruno Fieramosca, Nena Moonier, Clay White, Shane Flood, Wendy VanWingeren, Dean Wyncott, and Patti Pedergnana
- Bruce reminded the group to APS disable all stations besides 24-ID and 35-BM first thing tomorrow, by 0830 at the latest. Patti will take all beamlines global offline besides 24-ID and 35-BM, which should be done before 1000.
- Bruce was in an EA Assessment this morning and Sector Orientation was given by Sector 12 and Sector 35 beamline personnel.
- Bruce reminded the group to flush their area eyewash stations weekly.
- U5 Switchgear maintenance is this weekend. Next Saturday is the U3 and U4 Switchgear Maintenance, where there will be no cooling to the chilled water but the water will be flowing in the chilled water pipes.
- Dean is working on a WPC/JSA for the University of Rochester to install a laser target chamber at 35-IDC.
- 3-ID, 7-ID, and 33ID will cut into labyrinths to run LN2 lines. These stations will get the updated model ODH monitors.
- Patti has all FE CCWPs printed and ready to post.
- Patti reminded the group to please add all Experiment Floor CCWPs to the master list of CCWPs in the UES Share folder.
- The 9-ID mini-hutch splitting work begins tomorrow with Scheck removing the electrical. That will be followed by MOM and riggers moving the actual mini-hutches to the Sector 27 area to be stored until the upgrade is done at 2-ID. Steve Davey has a work plan for this job.
- Wendy will begin quarterly inventory checks of the APS MBA along with a DOE field inspection this week on Thursday.
- Nena reminded the group to close ESAFs, especially the proprietary experiments, as soon as possible once the shutdown begins.
- The biosafety cabinets will be certified this month.
- Sector 22 will be relocating a support table that will eventually hold a mirror. This work is being coordinated with S&A so a CCWP will be posted.